# Sharpness of Minima in Deep Matrix Factorization: Exact Expressions

### v4

@@ 6 Discussion
In this section, we thoroughly examine the claims made by Ding et al. (2024) and Ghosh et al. (2025) that flat minima coincide with Frobenius-norm balanced minima in depth-2 matrix factorization and then we discuss how misleading sharpness measures can be for the loss landscape analysis.
@@
Strict assumptions. First, we examine the latter work, where Ghosh et al. (2025) examined the learning dynamics of deep linear networks within the deep matrix factorization loss beyond the edge of stability. Their analysis relies on two strict assumptions, which the latter follows the first one. First, they introduce the singular vector stationary set, i.e., for any initialization of GD from this set, GD does not rotate the layers during training. In Proposition 4, they prove that the balanced initialization they considered in the paper, i.e., $\bm{W}_{L}(0)=\bm{W}_{L-1}(0)=\cdots=\bm{W}_{1}(0)=\alpha\bm{I}$ where $\alpha\in\mathbb{R}$, is a member of the singular vector stationary set. This is leveraged to decouple the dynamics of the singular vectors and singular values. Second, for this specific initialization, singular values remain balanced during training, which is not true for arbitrary initialization. Under these assumptions, their optimization objective becomes a deep scalar factorization problem in which the spectral norm equals the Frobenius norm. Therefore, they claim that Frobenius-norm balanced minima correspond to flat minima in deep matrix factorization problem. However, this is only valid from the perspective of GD that initialized at a specific set of points. Globally, we can find minimizers that are flat but not Frobenius-norm balanced. To see this, consider a depth-2 matrix factorization problem, i.e.,
$\min_{\bm{L},\bm{R}}\mathcal{L}(\bm{L},\bm{R})\quad\text{such that}\quad\mathcal{L}(\bm{L},\bm{R})=\norm{\bm{M}-\bm{L}\bm{R}^{\top}}_{F}^{2},$ (37)
@@
where $\bm{M}\in\mathbb{R}^{m\times n},\bm{L}\in\mathbb{R}^{m\times k}$ and $\bm{R}\in\mathbb{R}^{n\times k}$. Suppose $m=n=k=3$ and $\bm{M}=\text{diag}(11,2,1)$. We know that for any flat minimum $(\bm{L}^{\prime},\bm{R}^{\prime})$, $\lambda_{\max}(\nabla^{2}(\mathcal{L}(\bm{L}^{\prime},\bm{R}^{\prime})))=4\times\sigma_{\max}(\bm{M})$, which equals $44$ (Mulayoff & Michaeli, 2020). By definition, any minimum whose sharpness is equal to $44$ is a flat minimum. Let’s investigate a specific minimizer $(\bm{L}^{*},\bm{R}^{*})$ such that $\bm{L}^{*}=\text{diag}(\sqrt{11},\frac{2}{3},1)$ and $\bm{R}^{*}=\text{diag}(\sqrt{11},3,1)$. Using Corollary 25, $\lambda_{\max}(\nabla^{2}(\mathcal{L}(\bm{L}^{*},\bm{R}^{*})))=44$, which means that $(\bm{L}^{*},\bm{R}^{*})$ is a flat minimum. On the other hand, by definition of balancedness, i.e., $\bm{L}\bm{L}^{\top}=\bm{R}\bm{R}^{\top}$, $(\bm{L}^{*},\bm{R}^{*})$ is not a balanced minimizer. Therefore, any flat minimum is not necessarily a balanced minimizer.
@@
Sharpness measures are delusive. Ding et al. (2024) showed that flat minima recover the groundtruth matrix in low rank matrix recovery. Furthermore, they showed that norm-minimal, balanced, and flat solutions coincide in depth-$2$ matrix factorization. They measure sharpness of a minimum by using scaled trace. Under this measure, flat minima coincide with Frobenius-norm balanced minima; however, as we showed above, this is not necessarily true when the worst-case sharpness, i.e., maximum Hessian eigenvalue, is used as the sharpness measure. In fact, we showed that flat minima are spectral-norm balanced when the maximum Hessian eigenvalue is used as the sharpness measure. This means that a loss landscape analysis in general matrix factorization problems using a specific sharpness measure might not lead to the same inferences as those made by using a different metric.
@@
The necessity of an absolute sharpness measure. There is still no consensus in the literature regarding the definition of flatness. For instance, while several works define flat minima as global minimizers that minimize the maximum eigenvalue of the Hessian of the loss (Mulayoff & Michaeli, 2020; Liu et al., 2021; Marion & Chizat, 2024), others define them as global minimizers that minimize the trace of the Hessian (Dinh et al., 2017; Gatmiry et al., 2023). Even a scaled version of the Hessian trace designed to define flat minima in depth-2 matrix factorization problems (Ding et al., 2024), as we mentioned before. As we have shown in Section 5 and discussed above, different sharpness measures can lead to contradictory interpretations of balanced minima. Thus, the necessity of a robust sharpness measure that consolidates the interpretations from existing ones is a matter of urgency. Addressing this concern, Josz (2025) showed how minimizing the maximum Hessian eigenvalue over the solution set can disregard the higher-order variation near minimizers and result in misleading interpretations of flat minima. Therefore, Josz (2025) defined the flat minima of any smooth function $\mathcal{L}$ as the local minima of $\lambda_{\max}(\nabla^{2}\mathcal{L}(\bm{w}))$ under the constraint $\mathcal{L}(\bm{w})=\mathcal{L}(\bm{w}^{*})$ and demonstrated that this notion coincides with other notions of flatness in the depth-2 matrix factorization.
@@
(a) Normalized training error across iterations, i.e., $\mathcal{L}(\bm{w}_{k})/\lVert\bm{M}\rVert_{F}^{2}$.
(b) Normalized $\ell^{2}$ distance of $\bm{w}_{k}$ from the minimum $\bm{w}^{*}$, i.e, $\norm{\bm{w}_{k}-\bm{w}^{*}}_{2}^{2}/\norm{\bm{w}^{*}}_{2}^{2}$.
@@
Figure 1: GD dynamics with different step sizes, $\eta\geq 2/\lambda_{\max}(\nabla^{2}\mathcal{L}(\bm{w}^{*}))$, indicated by different colors, are initialized within a radius of $10^{-15}$ from the minimum in the direction of the Hessian eigenvector corresponding to the largest eigenvalue, for depth-22 matrix factorization, $\bm{M}=\bm{L}\bm{R}^{\top}$, of a random Gaussian matrix, where $\bm{L}\in\mathbb{R}^{10\times 20}$ and $\bm{R}\in\mathbb{R}^{20\times 20}$. The vector $\bm{v}_{1}$ denotes the eigenvector of the Hessian corresponding to the largest eigenvalue, while $\bm{v}_{N}$ denotes the eigenvector corresponding to the smallest eigenvalue. The value of $\lambda_{\max}(\nabla^{2}\mathcal{L}(\bm{w}^{*}))$ is computed using the closed-form expression derived in Corollary 25.
(a) Normalized training error across iterations, i.e., $\mathcal{L}(\bm{w}_{k})/m^{2}$.
(b) Normalized $\ell^{2}$ distance of $\bm{w}_{k}$ from the minimum $\bm{w}^{*}$.
(c) Trajectories of GD on the contour map of the loss landscape.
(d) Normalized training error across iterations, i.e., $\mathcal{L}(\bm{w}_{k})/m^{2}$.
(e) Normalized $\ell^{2}$ distance of $\bm{w}_{k}$ from the minimum $\bm{w}^{*}$.
(f) Trajectories of GD on the contour map of the loss landscape.
Figure 2: GD dynamics with different step sizes indicated by different colors are initialized within a radius of $10^{-12}$ from the minimum in the direction of the Hessian eigenvector corresponding to the largest eigenvalue, for a 15-layer overparameterized scalar factorization of a random scalar. The value of $\lambda_{\max}(\nabla^{2}\mathcal{L}(\bm{w}^{*}))$ is computed using the closed-form expression derived in Theorem 21.
@@ 7 Experiments
The dynamical stability analysis relies on how accurate the quadratic approximation of the loss function is in the $\delta$-neighborhood of $\bm{x}^{*}$ (Wu et al., 2018) (also see Definition 1). Therefore, we initialize GD in a $\delta$-neighborhood with $\delta$ on the order of $10^{-15}$ to $10^{-9}$ to ensure validity of the quadratic approximation. To observe escape phenomenon clearly, we choose the perturbation direction for the initial point to be the eigenvector of $\nabla^{2}\mathcal{L}(\bm{w}^{*})$ corresponding to the largest eigenvalue, so as to avoid choosing a direction that is orthogonal to the eigenspace corresponding to the eigenvalues larger than $2/\eta$.
@@
Note that, we could have also considered a random perturbation direction uniformly from $\mathbb{S}^{N-1}$. Mulayoff & Michaeli (2020) showed that for the deep matrix factorization problem (which has the same Hessian structure with deep linear networks), the Hessian is rank-deficient at all minima by at least the order of 1(1/L)1-(1/L). This means that at least 1(1/L)1-(1/L) of the eigenvalues are zero at a minimum. Thus, the probability of choosing a direction that is orthogonal to the eigenspace corresponding to eigenvalues larger than $2/\eta$ is 0. Therefore, random perturbations would lead to the escape phenomenon with probability $1$.
@@
On the other hand, for stable minima, it is important to choose the direction as the eigenvector of $\nabla^{2}\mathcal{L}(\bm{w}^{*})$ corresponding to the largest eigenvalue. The reasoning is the same. If you choose a direction that is not orthogonal to the eigenspace corresponding to the eigenvalues that are zero, then GD never converges to $\bm{w}^{*}$. This means that if we choose the perturbation direction randomly, then with probability 1, we choose a direction that is not orthogonal to the eigenspace corresponding to the eigenvalues that are zero. Therefore, for the experiment, it is convenient to choose perturbation direction as the eigenvector of $\nabla^{2}\mathcal{L}(\bm{w}^{*})$ corresponding to the largest eigenvalue.
@@
To measure the distance between the convergence point and the minimizer, we plot the normalized $\ell^{2}$-norm of $\bm{w}_{k}-\bm{w}^{*}$ at each iteration. Furthermore, as shown in Fig. 1 and Figs. 2(a)-2(c) if $\eta>2/\lambda_{\max}$, where $\lambda_{\max}:=\lambda_{\max}(\nabla^{2}\mathcal{L}(\bm{w}^{*}))$, GD always escapes from the minimum. On the other hand, if $\eta=2/\lambda_{\max}$ then GD converges as shown in Figs. 2(d)-2(f). A catapult in the training error indicates GD’s escape from the basin of a minimum, after which it eventually converges to another minimum as observed by Wu et al. (2018). For the methodology used to generate contour maps of the loss landscape near a minimum, see Appendix D.1, and for additional experiments, see Appendix 3.
@@ 8 Conclusion
In this paper, we derived an exact expression for the maximum eigenvalue of the Hessian of the squared-error loss for overparameterized deep matrix factorization problems at any minimizer. We also showed that this expression simplifies considerably for depth-2 matrix factorization and deep, overparameterized scalar factorization. To complement our theory, we conducted GD experiments that crucially rely on our exact expression of the sharpness to observe the escape phenomenon during training. Then, we showed how flat minima are spectral-norm balanced in depth-2 matrix factorization, and we showed that a minimizer of deep overparameterized scalar factorization loss is flat if and only if the product of spectral norms of left and right intermediate factors is constant across layers. Furthermore, we showed that a minimizer of deep matrix factorization loss is flat if the product of spectral norms of left and right intermediate factors is constant across layers. These results led to the following remarkable observation: flat minima are not necessarily Frobenius-norm balanced, which is contradictory to the claim made in (Ghosh et al., 2025). Finally, we discussed how delusive the sharpness measures can be for the loss landscape analysis. Therefore, there is an urgent need for future research to explore a robust sharpness measure that consolidates the interpretations from existing ones.
@@ A.1 Proof of Lemma 11
We can express the derivative of $\frac{\partial f(\bm{X})}{\partial X_{ij}}$ with respect to each entry of $\bm{X}$, using the limit definition of the derivative, as follows:
$\frac{\partial^{2}f(\bm{X})}{\partial X_{kl}\partial X_{ij}}=\frac{\partial}{\partial X_{kl}}\Bigl(\frac{\partial f(\bm{X})}{\partial X_{ij}}\Bigr)=\lim_{\Delta t\rightarrow 0}\frac{\partial f(\bm{X}+\Delta t\bm{e}_{i}\bm{e}_{j}^{\top})-\partial f(\bm{X})}{\partial X_{kl}\Delta t},\quad\forall(k,l)\in[K]\times[L]$ (38)
@@
$\lim_{\Delta h,\Delta t\rightarrow 0}\frac{f(\bm{X}+\Delta t\bm{e}_{i}\bm{e}_{j}^{\top}+\Delta h\bm{e}_{k}\bm{e}_{l}^{\top})-f(\bm{X}+\Delta t\bm{e}_{i}\bm{e}_{j}^{\top})-f(\bm{X}+\Delta h\bm{e}_{k}\bm{e}_{l}^{\top})+f(\bm{X})}{\Delta h\Delta t}.$ (39)
@@
$\frac{\partial^{2}f(\bm{X})}{\partial X_{kl}\partial X_{ij}}U_{ij}U_{kl}=\frac{\partial}{\partial X_{kl}}\Bigl(\frac{\partial f(\bm{X})}{\partial X_{ij}}U_{ij}\Bigr)U_{kl}=\lim_{\Delta t\rightarrow 0}\frac{\partial f(\bm{X}+\Delta tU_{ij}\bm{e}_{i}\bm{e}_{j}^{\top})-\partial f(\bm{X})}{\partial X_{kl}\Delta t}U_{kl}.$ (40)
@@
$\lim_{\Delta h,\Delta t\rightarrow 0}\frac{f(\bm{X}+\Delta tU_{ij}\bm{e}_{i}\bm{e}_{j}^{\top}+\Delta hU_{kl}\bm{e}_{k}\bm{e}_{l}^{\top})-f(\bm{X}+\Delta tU_{ij}\bm{e}_{i}\bm{e}_{j}^{\top})-f(\bm{X}+\Delta hU_{kl}\bm{e}_{k}\bm{e}_{l}^{\top})+f(\bm{X})}{\Delta h\Delta t}.$ (41)
which can be proved by substitution of variables in (40). In turn, second order differential due to any $\bm{U}\in\mathbb{R}^{K\times L}$ is
$D_{\bm{U}}^{2}f(\bm{X})=\sum_{i,j}\sum_{k,l}\frac{\partial^{2}f(\bm{X})}{\partial X_{kl}\partial X_{ij}}U_{ij}U_{kl}=\Bigl<\nabla\bigl<\nabla f(\bm{X}),\bm{U}\bigr>,\bm{U}\Bigr$>, (42)
@@
$\nabla f(\bm{X})=\begin{bmatrix}\frac{\partial f(\bm{X})}{\partial X_{11}}&\frac{\partial f(\bm{X})}{\partial X_{12}}&\dots&\frac{\partial f(\bm{X})}{\partial X_{1L}}\\ \frac{\partial f(\bm{X})}{\partial X_{21}}&\frac{\partial f(\bm{X})}{\partial X_{22}}&\dots&\frac{\partial f(\bm{X})}{\partial X_{2L}}\\ \vdots&\ddots&\dots&\vdots\\ \frac{\partial f(\bm{X})}{\partial X_{K1}}&\frac{\partial f(\bm{X})}{\partial X_{K2}}&\dots&\frac{\partial f(\bm{X})}{\partial X_{KL}}\\ \end{bmatrix}\in\mathbb{R}^{K\times L}.$ (43)
@@
$D_{\bm{U}}^{2}f(\bm{X})=$ $\sum_{k,l}\lim_{\Delta t\rightarrow 0}\frac{\partial f(\bm{X}+\Delta t\bm{U}\bm{e}_{i}\bm{e}_{j}^{\top})-\partial f(\bm{X})}{\partial X_{kl}\Delta t}U_{kl}$ (44)
$=$ $\lim_{\Delta t\rightarrow 0}\frac{f(\bm{X}+2\Delta t\bm{U})-2f(\bm{X}+\Delta t\bm{U})+f(\bm{X})}{\Delta t^{2}}$ (45)
$=$ $\frac{\partial^{2}}{\partial t^{2}}f(\bm{X}+t\bm{U})\Biggl|_{t=0}.$ (46)
A.2 Proof of Lemma 14
@@
We can rewrite (12) through using the definition of Frobenius inner product
$f(\bm{W}_{1},\cdots,\bm{W}_{L})=\Bigl<\bm{M}-\prod_{i=1}^{L}\bm{W}_{i},\bm{M}-\prod_{i=1}^{L}\bm{W}_{i}\Bigr$>. (47)
@@
$f(\bm{W}_{1}+t\bm{U}_{1},\cdots,\bm{W}_{L}+t\bm{U}_{L})=\Bigl<\bm{M}-\prod_{i=1}^{L}(\bm{W}_{i}+t\bm{U}_{i}),\bm{M}-\prod_{i=1}^{L}(\bm{W}_{i}+t\bm{U}_{i})\Bigr$>. (48)

### v5

@@ 5 Discussion
In this section, we thoroughly examine the claims made by Ding et al. (2024) and Ghosh et al. (2025) that flat minima coincide with Frobenius-norm balanced minima in depth-$2$ matrix factorization. We then discuss how sharpness measures can be misleading for loss landscape analysis.
5.1 Strict Assumptions
@@
First, we examine the latter work, where Ghosh et al. (2025) examined the learning dynamics of deep linear networks under the deep matrix factorization loss beyond the edge of stability. Their analysis relies on two strict assumptions, where the second follows from the first one. First, they introduce the singular vector stationary set, i.e., for any initialization of GD from this set, GD does not rotate the layers during training. In Proposition 4, they prove that the Frobenius-norm balanced initialization they considered in the paper, i.e., $\mathbf{W}_{L}(0)=\mathbf{W}_{L-1}(0)=\cdots=\mathbf{W}_{1}(0)=\alpha\mathbf{I}$ where $\alpha\in\mathbb{R}$, is a member of the singular vector stationary set. This is leveraged to decouple the dynamics of the singular vectors and singular values. Second, for this specific initialization, singular values remain balanced during training, which is not true for arbitrary initialization. Under these assumptions, their optimization objective becomes a deep scalar factorization problem in which the spectral norm equals the Frobenius norm. Therefore, they claim that Frobenius-norm balanced minima correspond to flat minima in deep matrix factorization problem. However, this is only valid from the perspective of GD initialized at a specific set of points. Globally, we can find minimizers that are flat but not Frobenius-norm balanced. To see this, consider a depth-$2$ matrix factorization problem, i.e.,
$\min_{\mathbf{L},\mathbf{R}}\mathcal{L}(\mathbf{L},\mathbf{R})\>\>\text{where}\>\>\mathcal{L}(\mathbf{L},\mathbf{R})=\norm{\mathbf{M}-\mathbf{L}\mathbf{R}^{\top}}_{F}^{2},$ (40)
@@
where $\mathbf{M}\in\mathbb{R}^{m\times n},\mathbf{L}\in\mathbb{R}^{m\times k}$ and $\mathbf{R}\in\mathbb{R}^{n\times k}$. Suppose $m=n=k=3$ and $\mathbf{M}=\text{diag}(11,2,1)$. We know that for any flat minimum $(\mathbf{L}^{\prime},\mathbf{R}^{\prime})$, $\lambda_{\max}(\nabla^{2}(\mathcal{L}(\mathbf{L}^{\prime},\mathbf{R}^{\prime})))=4\times\sigma_{\max}(\mathbf{M})$, which equals $44$. By definition, any minimum whose sharpness is equal to $44$ is a flat minimum. Let’s investigate a specific minimizer $(\mathbf{L}^{*},\mathbf{R}^{*})$ such that $\mathbf{L}^{*}=\text{diag}(\sqrt{11},\frac{2}{3},1)$ and $\mathbf{R}^{*}=\text{diag}(\sqrt{11},3,1)$. Using Corollary 28, $\lambda_{\max}(\nabla^{2}(\mathcal{L}(\mathbf{L}^{*},\mathbf{R}^{*})))=44$, which means that $(\mathbf{L}^{*},\mathbf{R}^{*})$ is a flat minimum. On the other hand, by definition of Frobenius-norm balancedness, i.e., $\mathbf{L}\mathbf{L}^{\top}=\mathbf{R}\mathbf{R}^{\top}$, $(\mathbf{L}^{*},\mathbf{R}^{*})$ is not Frobenius-norm balanced. Therefore, a flat minimum is not necessarily Frobenius-norm balanced.
5.2 Sharpness Measures Are Delusive
@@
Ding et al. (2024) showed that flat minima recover the ground truth matrix in low-rank matrix recovery. Furthermore, they showed that norm-minimal, Frobenius-balanced and flat solutions coincide in depth-$2$ matrix factorization (Ding et al., 2024, Lemma 2.2). They measured the sharpness of a minimum using scaled trace. Under this measure, flat minima coincide with Frobenius-norm balanced minima; however, as we showed above, this is not necessarily true when the worst-case sharpness, i.e., maximum Hessian eigenvalue, is used as the sharpness measure. In fact, we showed that flat minima are spectral-norm balanced when the maximum Hessian eigenvalue is used as the sharpness measure (Corollary 36). This means that a loss landscape analysis in general matrix factorization/two-layer linear neural network training problems using a specific sharpness measure might not lead to the same inferences as those made by using a different metric.
5.3 The Necessity of an Absolute Sharpness Measure
@@
There is still no consensus in the literature regarding the definition of flatness. For instance, while several works define flat minima as global minimizers that minimize the maximum eigenvalue of the Hessian of the loss (Mulayoff and Michaeli, 2020; Liu et al., 2021; Marion and Chizat, 2024), others define them as global minimizers that minimize the trace of the Hessian (Dinh et al., 2017; Gatmiry et al., 2023). Even a scaled version of the Hessian trace designed to define flat minima in depth-$2$ matrix factorization problems (Ding et al., 2024), as we mentioned before. As we have shown in Section 4 and discussed above, different sharpness measures can lead to contradictory interpretations of balanced minima. Thus, the necessity of a robust sharpness measure that consolidates the interpretations from existing ones is a matter of urgency. Addressing this concern, Josz (2025) showed how minimizing the maximum Hessian eigenvalue over the solution set can disregard the higher-order variation near minimizers and result in misleading interpretations of flat minima. Therefore, Josz (2025) defined the flat minima of any smooth function $\mathcal{L}$ as the local minima of $\lambda_{\max}(\nabla^{2}\mathcal{L}(\mathbf{w}))$ under the constraint $\mathcal{L}(\mathbf{w})=\mathcal{L}(\mathbf{w}^{*})$ and demonstrated that this notion coincides with other notions of flatness in the depth-$2$ matrix factorization.
@@
(a) Normalized training error across iterations, i.e., $\mathcal{L}(\mathbf{w}_{k})/\lVert\mathbf{M}\rVert_{F}^{2}$.
(b) Normalized $\ell^{2}$ distance of $\mathbf{w}_{k}$ from the minimum $\mathbf{w}^{*}$, i.e., $\norm{\mathbf{w}_{k}-\mathbf{w}^{*}}_{2}^{2}/\norm{\mathbf{w}^{*}}_{2}^{2}$.
@@
Figure 1: GD dynamics with different step sizes indicated by different colors are initialized within a radius of $10^{-15}$ from the minimum in the direction of the Hessian eigenvector that corresponds to the maximum eigenvalue for the depth-$2$ matrix factorization of a random Gaussian matrix, where $\mathbf{L}\in\mathbb{R}^{10\times 20}$ and $\mathbf{R}\in\mathbb{R}^{20\times 20}$. The vector $\mathbf{v}_{1}$ denotes the eigenvector of the Hessian that corresponds to the maximum eigenvalue, while $\mathbf{v}_{N}$ denotes the eigenvector that corresponds to the smallest eigenvalue. The value of $\lambda_{\max}(\nabla^{2}\mathcal{L}(\mathbf{w}^{*}))$ is computed using the closed-form expression derived in Corollary 28. Darker regions indicate higher loss, whereas brighter regions indicate lower loss.
(a) Trajectories of GD initialized at $\mathbf{w}_{0}$ with step sizes $\geq 2/{\lambda_{\max}}$ are depicted by colored lines, and their corresponding convergence points are marked by colored $\star$ symbols.
(b) Trajectories of GD initialized at $\mathbf{w}_{0}$ with step sizes $2/{\lambda_{\max}}$ are depicted by colored lines, and their corresponding convergence points are marked by colored $\star$ symbols.
Figure 2: Contour map of the loss landscape around a minimizer of a 15-layer overparameterized scalar factorization problem. GD with different step sizes indicated by different colors is initialized within a radius of $10^{-9}$ from the minimum in the direction of the Hessian eigenvector that corresponds to the maximum eigenvalue. The vector $\mathbf{v}_{1}$ denotes the eigenvector of the Hessian that corresponds to the maximum eigenvalue, while $\mathbf{v}_{N}$ denotes the eigenvector that corresponds to the smallest eigenvalue. The value of $\lambda_{\max}(\nabla^{2}\mathcal{L}(\mathbf{w}^{*}))$ is computed using the closed-form expression derived in Corollary 3.2. Darker regions indicate higher loss, whereas brighter regions indicate lower loss.
@@ 6 Experiments
The dynamical stability analysis relies on how accurate the quadratic approximation of the loss function is in the $\delta$-neighborhood of $\mathbf{w}^{*}$ (Wu et al., 2018) (also see Definition 5). For loss functions whose Hessian is degenerate at a minimizer, this analysis fails. In contrast, the squared-error loss ensures that the Hessian at any minimizer is not degenerate/does not vanish (also see Mulayoff and Michaeli (2020, Lemma 3)). Therefore, we can initialize GD in a $\delta$-neighborhood with $\delta$ of order $10^{-15}$ to $10^{-9}$ so that the linearized GD dynamics approximate the true GD dynamics well. To observe escape phenomenon clearly, we choose the perturbation direction for the initial point to be the eigenvector of $\nabla^{2}\mathcal{L}(\mathbf{w}^{*})$ corresponding to the largest eigenvalue, so as to avoid choosing a direction that is orthogonal to the eigenspace corresponding to the eigenvalues larger than $2/\eta$.
@@
Note that we could have also considered a random perturbation direction sampled uniformly from $\mathbb{S}^{N-1}$. Mulayoff and Michaeli (2020) showed that for the deep matrix factorization problem (which has the same Hessian structure with deep linear networks with a full rank data matrix), the Hessian is rank-deficient at all minima by at least the order of 1(1/L)1-(1/L). This means that at least 1(1/L)1-(1/L) of the eigenvalues are zero at a minimum. Thus, the probability of choosing a direction that is orthogonal to the eigenspace corresponding to eigenvalues larger than $2/\eta$ is 0. Therefore, random perturbations would lead to the escape phenomenon with probability $1$.
@@
On the other hand, for stable minima, it is important to choose the direction as the eigenvector of $\nabla^{2}\mathcal{L}(\mathbf{w}^{*})$ that corresponds to the largest eigenvalue. The reasoning is the same. If you choose a direction that is not orthogonal to the eigenspace corresponding to the eigenvalues that are zero, then GD never converges to $\mathbf{w}^{*}$. This means that if we choose the perturbation direction randomly, then with probability $1$, we choose a direction that is not orthogonal to the eigenspace corresponding to the eigenvalues that are zero. Therefore, for the experiment, it is convenient to choose the perturbation direction as the eigenvector of $\nabla^{2}\mathcal{L}(\mathbf{w}^{*})$ corresponding to the largest eigenvalue. Note that the convergence we refer to here is the convergence to a specific global minimizer, not to an arbitrary global minimizer.
@@
To measure the distance between the convergence point and the minimizer, we plot the normalized $\ell^{2}$-norm of $\mathbf{w}_{k}-\mathbf{w}^{*}$ at each iteration as shown in Fig. 1. Furthermore, if $\eta>2/\lambda_{\max}$, where $\lambda_{\max}:=\lambda_{\max}(\nabla^{2}\mathcal{L}(\mathbf{w}^{*}))$, GD always escapes from the minimum as shown in Fig. 1 and Fig. 2. On the other hand, as shown in Fig. 2, if $\eta\leq 2/\lambda_{\max}$, then GD converges. A catapult in the training error shown in Fig. 1 indicates GD’s escape from the basin of a minimum, after which it eventually converges to another minimum unless the step size is very large (Marion and Chizat, 2024). To the best of our knowledge, this is the first empirical characterization of an escape phenomenon during gradient-based training near a minimizer of a deep matrix factorization problem. For the methodology used to generate contour maps of the loss landscape near a minimum, see Appendix D.1, and for additional experiments, see Appendix D.2.
@@ 7 Conclusion
In this paper, we derived an exact expression for the maximum eigenvalue of the Hessian of the squared-error loss in overparameterized deep matrix factorization at any minimizer, and showed that it simplifies substantially in the depth-$2$ matrix factorization and deep overparameterized scalar factorization settings. We further established that flat minima are spectral-norm balanced in depth-$2$ matrix factorization, and that a minimizer of deep overparameterized scalar factorization is flat if and only if the product of the spectral norms of the left and right intermediate factors is constant across layers; the same condition is sufficient for flatness in deep matrix factorization. To complement our theory, we conducted GD experiments that crucially rely on our exact expression of the sharpness to observe the escape phenomenon during training. Finally, we discussed that sharpness measures can be misleading for loss landscape analysis, highlighting the need for robust sharpness notions that reconcile existing interpretations.
@@ A.1 Proof of Lemma 14
We can express the derivative of $\frac{\partial f(\mathbf{X})}{\partial X_{ij}}$ with respect to each entry of $\mathbf{X}$, using the limit definition of the derivative, as follows:
$\frac{\partial^{2}f(\mathbf{X})}{\partial X_{kl}\partial X_{ij}}=\frac{\partial}{\partial X_{kl}}\Bigl(\frac{\partial f(\mathbf{X})}{\partial X_{ij}}\Bigr)=\lim_{\Delta t\rightarrow 0}\frac{\partial f(\mathbf{X}+\Delta t\mathbf{e}_{i}\mathbf{e}_{j}^{\top})-\partial f(\mathbf{X})}{\partial X_{kl}\Delta t},\quad\forall(k,l)\in[K]\times[L]$ (41)
@@
$\lim_{\Delta h,\Delta t\rightarrow 0}\frac{f(\mathbf{X}+\Delta t\mathbf{e}_{i}\mathbf{e}_{j}^{\top}+\Delta h\mathbf{e}_{k}\mathbf{e}_{l}^{\top})-f(\mathbf{X}+\Delta t\mathbf{e}_{i}\mathbf{e}_{j}^{\top})-f(\mathbf{X}+\Delta h\mathbf{e}_{k}\mathbf{e}_{l}^{\top})+f(\mathbf{X})}{\Delta h\Delta t}.$ (42)
@@
$\frac{\partial^{2}f(\mathbf{X})}{\partial X_{kl}\partial X_{ij}}U_{ij}U_{kl}=\frac{\partial}{\partial X_{kl}}\Bigl(\frac{\partial f(\mathbf{X})}{\partial X_{ij}}U_{ij}\Bigr)U_{kl}=\lim_{\Delta t\rightarrow 0}\frac{\partial f(\mathbf{X}+\Delta tU_{ij}\mathbf{e}_{i}\mathbf{e}_{j}^{\top})-\partial f(\mathbf{X})}{\partial X_{kl}\Delta t}U_{kl}.$ (43)
@@
$\lim_{\Delta h,\Delta t\rightarrow 0}\frac{f(\mathbf{X}+\Delta tU_{ij}\mathbf{e}_{i}\mathbf{e}_{j}^{\top}+\Delta hU_{kl}\mathbf{e}_{k}\mathbf{e}_{l}^{\top})-f(\mathbf{X}+\Delta tU_{ij}\mathbf{e}_{i}\mathbf{e}_{j}^{\top})-f(\mathbf{X}+\Delta hU_{kl}\mathbf{e}_{k}\mathbf{e}_{l}^{\top})+f(\mathbf{X})}{\Delta h\Delta t}.$ (44)
which can be proved by substitution of variables in (43). In turn, second order differential due to any $\mathbf{U}\in\mathbb{R}^{K\times L}$ is
$D_{\mathbf{U}}^{2}f(\mathbf{X})=\sum_{i,j}\sum_{k,l}\frac{\partial^{2}f(\mathbf{X})}{\partial X_{kl}\partial X_{ij}}U_{ij}U_{kl}=\Bigl<\nabla\bigl<\nabla f(\mathbf{X}),\mathbf{U}\bigr>,\mathbf{U}\Bigr$>, (45)
@@
$\nabla f(\mathbf{X})=\begin{bmatrix}\frac{\partial f(\mathbf{X})}{\partial X_{11}}&\frac{\partial f(\mathbf{X})}{\partial X_{12}}&\dots&\frac{\partial f(\mathbf{X})}{\partial X_{1L}}\\ \frac{\partial f(\mathbf{X})}{\partial X_{21}}&\frac{\partial f(\mathbf{X})}{\partial X_{22}}&\dots&\frac{\partial f(\mathbf{X})}{\partial X_{2L}}\\ \vdots&\ddots&\dots&\vdots\\ \frac{\partial f(\mathbf{X})}{\partial X_{K1}}&\frac{\partial f(\mathbf{X})}{\partial X_{K2}}&\dots&\frac{\partial f(\mathbf{X})}{\partial X_{KL}}\\ \end{bmatrix}\in\mathbb{R}^{K\times L}.$ (46)
@@
$D_{\mathbf{U}}^{2}f(\mathbf{X})=$ $\sum_{k,l}\lim_{\Delta t\rightarrow 0}\frac{\partial f(\mathbf{X}+\Delta t\mathbf{U}\mathbf{e}_{i}\mathbf{e}_{j}^{\top})-\partial f(\mathbf{X})}{\partial X_{kl}\Delta t}U_{kl}$ (47)
$=$ $\lim_{\Delta t\rightarrow 0}\frac{f(\mathbf{X}+2\Delta t\mathbf{U})-2f(\mathbf{X}+\Delta t\mathbf{U})+f(\mathbf{X})}{\Delta t^{2}}$ (48)
$=$ $\frac{\partial^{2}}{\partial t^{2}}f(\mathbf{X}+t\mathbf{U})\Biggl|_{t=0}.$ (49)
A.2 Proof of Lemma 2.2
@@
We can rewrite (15) by using the definition of Frobenius inner product
$f(\mathbf{W}_{1},\cdots,\mathbf{W}_{L})=\Bigl<\mathbf{M}-\prod_{i=1}^{L}\mathbf{W}_{i},\mathbf{M}-\prod_{i=1}^{L}\mathbf{W}_{i}\Bigr$>. (50)
@@
$f(\mathbf{W}_{1}+t\mathbf{U}_{1},\cdots,\mathbf{W}_{L}+t\mathbf{U}_{L})=\Bigl<\mathbf{M}-\prod_{i=1}^{L}(\mathbf{W}_{i}+t\mathbf{U}_{i}),\mathbf{M}-\prod_{i=1}^{L}(\mathbf{W}_{i}+t\mathbf{U}_{i})\Bigr$>. (51)